# Bridging Local Details and Global Context in Text-Attributed Graphs

### v1

@@ Abstract
Representation learning on text-attributed graphs (TAGs) is vital for real-world applications, as they combine semantic textual and contextual structural information. Research in this field generally consist of two main perspectives: local-level encoding and global-level aggregating, respectively refer to textual node information unification (e.g., using Language Models) and structure-augmented modeling (e.g., using Graph Neural Networks). Most existing works focus on combining different information levels but overlook the interconnections, i.e., the contextual textual information among nodes, which provides semantic insights to bridge local and global levels. In this paper, we propose GraphBridge, a multi-granularity integration framework that bridges local and global perspectives by leveraging contextual textual information, enhancing fine-grained understanding of TAGs. Besides, to tackle scalability and efficiency challenges, we introduce a graph-aware token reduction module. Extensive experiments across various models and datasets show that our method achieves state-of-the-art performance, while our graph-aware token reduction module significantly enhances efficiency and solves scalability issues.
@@
Yaoke Wang $1$These authors contributed equally to this work., Yun Zhu1111These authors contributed equally to this work., Wenqiao Zhang1, Yueting Zhuang1, Yunfei Li2, Siliang Tang1222Corresponding Author 1Zhejiang University, 2Ant Group {wangyaoke, zhuyun_dcd, wenqiaozhang, yzhuang}@zju.edu.cn qixiu.lyf@antgroup.com, siliang@zju.edu.cn
@@ 1 Introduction
Text-Attributed Graphs (TAGs), characterized by the association of nodes with text attributes Yang et al. (2021), are prevalent in diverse real-world contexts. In TAGs, nodes represent entities with textual information and edges capture relationships between entities, e.g., social graphs where each user is accompanied by a textual description and paper citation graphs where textual content is linked to each respective paper. These relationships yield specialized and crucial insights that are fundamental for our understanding, thereby facilitating the resolution of subsequent tasks. The utilization of TAGs empowers us to unlock new discoveries across various domains, including graph learning Zhang et al. (2024) and information retrieval Seo et al. (2024).
@@ 3.2 Graph-Aware Token Reduction
where $\text{Score}_{i}\in\mathbb{R}^{1\times k}$ is the importance score for each token in node $i$, and $\mathcal{N}_{i}$ denotes the neighboring nodes of $i$. The importance score is calculated by a trainable graph-enhanced attention module, as depicted in Figure 2.
@@ 3.3 Multi-Granularity Integration
As discussed in Section 1, the integration of local-global perspectives is essential for representation learning in TAGs Yan et al. (2023). However, previous methods Duan et al. (2023); He et al. (2023) often fail to adequately address the interconnection between these two perspectives. In this work, we propose an innovative multi-granularity integration framework that bridges local and global perspectives by considering contextual textual semantics among nodes, as illustrated in Figure 2. Intuitively, the contextual textual semantics among nodes can offer supplementary information for a given node. For example, within a citation network, the textual content of neighboring nodes might include key terms or concepts pertinent to the target node.
@@ 4 Experiments
In this section, we first introduce the used datasets in Section 4.1. We then detail the baseline methods and experimental setup in Sections 4.2 and 4.3, respectively. Experiments are presented to evaluate our proposed method in Section 4.4. We further investigate the use of a causal large language model as the backbone in Section 4.5. Finally, we provide an analysis of hyper-parameters, assess scalability and efficiency, and conduct an ablation study.
@@ 4.5 Enhanced with Large Language Models
Our method, as demonstrated in Section 4.4, proves effective with small and medium discriminative LMs like RoBERTa-base and RoBERTa-large. Furthermore, we have expanded our method to incorporate causal Large Language Models (LLMs), which have shown significant capabilities across various natural language tasks Achiam et al. (2023). Table 3 presents the results obtained using LLaMA2-7B Touvron et al. (2023) as the LLM backbone. Our method outperforms both LM-based method (i.e., LLaMA2) and integration methods (i.e., SimTeG, ENGINE) that utilize LLM. This demonstrates the effectiveness of our method with the LLM backbone, highlighting the importance of bridging the local and global perspectives.
Methods Cora WikiCS Ele-Photo
LLaMA2 82.80 ± 1.37 80.82 ± 0.48 72.06 ± 0.10
$\text{SimTeG}_{(\text{ LLaMA2})}$ 92.84 ± 0.13 82.55 ± 0.51 82.05 ± 0.17
$\text{ENGINE}_{(\text{ LLaMA2})}$ 91.48 ± 0.32 81.56 ± 0.97 83.75 ± 0.08
$\text{Ours}_{(\text{ LLaMA2})}$ 93.65 ± 0.44 84.18 ± 0.68 84.35 ± 0.12
Figure 4: Sensitive analysis of the number of walk steps.
@@ 64 OOM --
The regularization term $\beta$.
4.7 Scalability and Efficiency Analysis
@@ Appendix B Baselines
  • GLEM Zhao et al. (2022) is an effective framework that fuses language models and GNNs in the training phase through a variational EM framework. We use the official source codehttps://github.com/AndyJZhao/GLEM to reproduce its results.
  • TAPE He et al. (2023) utilizes Large Language Models like ChatGPT Achiam et al. (2023) to generate pseudo labels and explanations for textual nodes, which are then used to fine-tune Pre-trained Language Models alongside the original texts. We reproduced its results using the official source codehttps://github.com/XiaoxinHe/TAPE .
  • SimTeG Duan et al. (2023) employs a cascading structure specifically designed for textual graphs, utilizing a two-stage training paradigm. Initially, it fine-tunes language models and subsequently trains GNNs. We conducted experiments using the official source code§§§https://github.com/vermouthdky/SimTeG.
  • ENGINE Zhu et al. (2024) is an efficient fine-tuning and inference framework for text-attributed graphs. It co-trains large language models and GNNs using a ladder-side approach, optimizing both memory and time efficiency. For inference, ENGINE utilizes an early exit strategy to further accelerate. We reproduce its results using the official source codehttps://github.com/ZhuYun97/ENGINE.
@@ Appendix C Implementation Details
We train only the cross-attention module and the classifier, keeping the encoding language models frozen. Each dataset undergoes 100 training epochs, with an early stopping patience of 10 epochs. The learning rate is explored within {1e-3, 5e-4, 1e-4}, and the regularization term $\beta$ is set to 0.1.
@@
Initially, we construct the sequence $Q$ by sampling adjacent nodes using a random walk with restart sampler $\Gamma$ Zhu et al. (2022, 2024, 2023). The number of walk steps for sampling varies among {8, 16, 32}. Training epochs for the language models are adapted according to the dataset sizes: {4, 6, 8} for small datasets (e.g., Cora, WikiCS, CiteSeer), {4, 6} for medium datasets (e.g., ArXiv-2023, Ele-Photo, OGBN-Products), and {4} for the large-scale dataset (OGBN-ArXiv). We employ AdamW optimizers. The learning rate is explored within {1e-4, 5e-5, 1e-5} for full parameter fine-tuning of RoBERTa-basehttps://huggingface.co/FacebookAI/roberta-base, and {1e-3, 5e-4, 1e-4} for tuning RoBERTa-large******https://huggingface.co/FacebookAI/roberta-large using LoRA Hu et al. (2021) with a rank of 8. For the large language model LLaMA2-7B††††††https://huggingface.co/meta-llama/Llama-2-7b, as outlined in Table 3, we use LoRA with a rank of 4 and a learning rate within {5e-4, 1e-4, 5e-5}.

### v2

@@ Abstract
Representation learning on text-attributed graphs (TAGs) is vital for real-world applications, as they combine semantic textual and contextual structural information. Research in this field generally consist of two main perspectives: local-level encoding and global-level aggregating, respectively refer to textual node information unification (e.g., using Language Models) and structure-augmented modeling (e.g., using Graph Neural Networks). Most existing works focus on combining different information levels but overlook the interconnections, i.e., the contextual textual information among nodes, which provides semantic insights to bridge local and global levels. In this paper, we propose GraphBridge, a multi-granularity integration framework that bridges local and global perspectives by leveraging contextual textual information, enhancing fine-grained understanding of TAGs. Besides, to tackle scalability and efficiency challenges, we introduce a graph-aware token reduction module. Extensive experiments across various models and datasets show that our method achieves state-of-the-art performance, while our graph-aware token reduction module significantly enhances efficiency and solves scalability issues. Codes are available at https://github.com/wykk00/GraphBridge
@@
Yaoke Wang $1$These authors contributed equally to this work., Yun Zhu1111These authors contributed equally to this work., Wenqiao Zhang1222Corresponding Author, Yueting Zhuang1, Yunfei Li2, Siliang Tang1 1Zhejiang University, 2Ant Group {wangyaoke, zhuyun_dcd, wenqiaozhang, yzhuang}@zju.edu.cn qixiu.lyf@antgroup.com, siliang@zju.edu.cn
@@ 1 Introduction
Text-Attributed Graphs (TAGs), characterized by the association of nodes with text attributes Yang et al. (2021), are prevalent in diverse real-world contexts. In TAGs, nodes represent entities with textual information and edges capture relationships between entities, e.g., social graphs where each user is accompanied by a textual description and paper citation graphs where textual content is linked to each respective paper. These relationships yield specialized and crucial insights that are fundamental for our understanding, thereby facilitating the resolution of subsequent tasks. The utilization of TAGs empowers us to unlock new discoveries across various domains, including graph learning Zhang et al. (2024a) and information retrieval Seo et al. (2024).
@@ 3.2 Graph-Aware Token Reduction
where $\text{Score}_{i}\in\mathbb{R}^{1\times k}$ is the importance score for each token in node $i$, and $\mathcal{N}_{i}$ denotes the neighboring nodes of $i$. The importance score is calculated by a trainable graph-enhanced attention module, as depicted in Figure 2 (Stage 1).
@@ 3.3 Multi-Granularity Integration
As discussed in Section 1, the integration of local-global perspectives is essential for representation learning in TAGs Yan et al. (2023). However, previous methods Duan et al. (2023); He et al. (2023) often fail to adequately address the interconnection between these two perspectives. In this work, we propose an innovative multi-granularity integration framework that bridges local and global perspectives by considering contextual textual semantics among nodes, as illustrated in Figure 2 (Stage 2). Intuitively, the contextual textual semantics among nodes can offer supplementary information for a given node. For example, within a citation network, the textual content of neighboring nodes might include key terms or concepts pertinent to the target node.
@@ 4 Experiments
In this section, we first introduce the used datasets in Section 4.1. We then detail the baseline methods and experimental setup in Sections 4.2 and 4.3, respectively. Experiments are presented to evaluate our proposed method in Section 4.4. We further investigate the use of a causal large language model as the backbone in Section 4.5. Finally, we provide an analysis of hyper-parameters, assess scalability and efficiency, and conduct an ablation study. Furthermore, additional experiments (e.g., link prediction) are in Appendix.
@@ 4.1 Datasets
Methods Cora WikiCS Ele-Photo
LLaMA2 82.80 ± 1.37 80.82 ± 0.48 72.06 ± 0.10
$\text{SimTeG}_{(\text{ LLaMA2})}$ 92.84 ± 0.13 82.55 ± 0.51 82.05 ± 0.17
$\text{ENGINE}_{(\text{ LLaMA2})}$ 91.48 ± 0.32 81.56 ± 0.97 83.75 ± 0.08
$\text{Ours}_{(\text{ LLaMA2})}$ 93.65 ± 0.44 84.18 ± 0.68 84.35 ± 0.12
@@ 4.5 Enhanced with Large Language Models
Our method, as demonstrated in Section 4.4, proves effective with small and medium discriminative LMs like RoBERTa-base and RoBERTa-large. Furthermore, we have expanded our method to incorporate causal Large Language Models (LLMs), which have shown significant capabilities across various natural language tasks Achiam et al. (2023); Zhang et al. (2024b, c). Table 3 presents the results obtained using LLaMA2-7B Touvron et al. (2023) as the LLM backbone. Our method outperforms both LM-based method (i.e., LLaMA2) and integration methods (i.e., SimTeG, ENGINE) that utilize LLM. This demonstrates the effectiveness of our method with the LLM backbone, highlighting the importance of bridging the local and global perspectives.
@@ 4.6 Sensitive Analysis
Figure 4: Sensitive analysis of the number of walk steps.
@@
The regularization term $\beta$.
4.7 Scalability and Efficiency Analysis
@@ 6 Limitations
7 Acknowledgements
This work has been supported in part by the Key Research and Development Projects in Zhejiang Province (No. 2024C01106), the NSFC (No. 62272411), the National Key Research and Development Project of China (2018AAA0101900), and Ant Group.
@@ Appendix B Baselines
  • Previous methods for TAGs: GLEM Zhao et al. (2022) is an effective framework that fuses language models and GNNs in the training phase through a variational EM framework. We use the official source codehttps://github.com/AndyJZhao/GLEM to reproduce its results. TAPE He et al. (2023) utilizes Large Language Models like ChatGPT Achiam et al. (2023) to generate pseudo labels and explanations for textual nodes, which are then used to fine-tune Pre-trained Language Models alongside the original texts. We reproduced its results using the official source codehttps://github.com/XiaoxinHe/TAPE . SimTeG Duan et al. (2023) employs a cascading structure specifically designed for textual graphs, utilizing a two-stage training paradigm. Initially, it fine-tunes language models and subsequently trains GNNs. We conducted experiments using the official source code§§§https://github.com/vermouthdky/SimTeG. ENGINE Zhu et al. (2024) is an efficient fine-tuning and inference framework for text-attributed graphs. It co-trains large language models and GNNs using a ladder-side approach, optimizing both memory and time efficiency. For inference, ENGINE utilizes an early exit strategy to further accelerate. We reproduce its results using the official source codehttps://github.com/ZhuYun97/ENGINE.
@@ Appendix C Implementation Details
We train only the cross-attention module and the classifier, keeping the encoding language models frozen Zhang et al. (2019, 2022). Each dataset undergoes 100 training epochs, with an early stopping patience of 10 epochs. The learning rate is explored within {1e-3, 5e-4, 1e-4}, and the regularization term $\beta$ is set to 0.1.
@@
Initially, we construct the sequence $Q$ by sampling adjacent nodes using a random walk with restart sampler $\Gamma$ Zhu et al. (2022, 2024). The number of walk steps for sampling varies among {8, 16, 32}. Training epochs for the language models are adapted according to the dataset sizes: {4, 6, 8} for small datasets (e.g., Cora, WikiCS, CiteSeer), {4, 6} for medium datasets (e.g., ArXiv-2023, Ele-Photo, OGBN-Products), and {4} for the large-scale dataset (OGBN-ArXiv). We employ AdamW optimizers. The learning rate is explored within {1e-4, 5e-5, 1e-5} for full parameter fine-tuning of RoBERTa-basehttps://huggingface.co/FacebookAI/roberta-base, and {1e-3, 5e-4, 1e-4} for tuning RoBERTa-large******https://huggingface.co/FacebookAI/roberta-large using LoRA Hu et al. (2021) with a rank of 8. For the large language model LLaMA2-7B††††††https://huggingface.co/meta-llama/Llama-2-7b, as outlined in Table 3, we use LoRA with a rank of 4 and a learning rate within {5e-4, 1e-4, 5e-5}.
@@ Appendix D Alternative LMs as Backbones for Token Reduction
Appendix E Link Prediction
Methods Cora CiteSeer ArXiv-2023
SAGE 86.20 ± 0.96 79.69 ± 1.11 94.90 ± 0.24
RoBERTa-base 83.68 ± 0.61 85.10 ± 1.23 92.59 ± 0.55
SimTeG 87.24 ± 0.47 86.33 ± 1.05 97.01 ± 0.64
$\text{Ours}_{(\text{base})}$ 88.87 ± 0.81 88.01 ± 1.20 97.33 ± 0.48
Table 6: Experimental results of link prediction. The mean AUC with standard deviation across 5 runs is reported.
Our method is not limited to node classification tasks, it can also be applied to other representation learning tasks of TAGs (e.g., link prediction). In this section, we conduct experiments on link prediction task. Specifically, we split the existing edges into train:val:test=0.6:0.2:0.2 for all datasets. The AUC score is served as the metric. According to Table 6, GraphBridge still outperforms other baselines by a significant margin. This highlights the effectiveness of its representation learning on TAGs and demonstrates the potential of GraphBridge for application to other-level tasks.